A monopolist firm faces a demand with constant elasticity of negative 1.8. It has a constant marginal cost of ​$15 per unit and sets a price to maximize profit. If marginal cost should increase by 20 ​percent, would the price charged also rise by 20 ​percent?

A. Yes. Since the price elasticity of demand is​ constant, Upper P equals 1.8 MC. ​Thus, if MC increases by 20 ​percent, price also increases by 20 percent.

B. Yes. Since the price elasticity of demand is​ constant, Upper P equals 2.25 MC. ​Thus, if MC increases by 20 ​percent, price also increases by 20 percent.

C. No. Since the demand curve is downward​ sloping, a 20 percent increase in MC will cause the price to increase by more than 20 percent.

D. No. Since the demand curve is downward​ sloping, a 20 percent increase in MC will cause the price to increase by less than 20 percent.
Answer #1

Calculate the old profit maximizing price as follows: P-MC_1 Р Е. P-15_1 P 1.8 P-15=P 1.8 2P-27 =P 2P - P = 27 p* = 27 Thus,P-MC Р P-18 1 Е. 1 P 1.8 P-18= 1.8 2P-32.4 = P 2P - P = 32.4 p* = $32.4 Thus, price is $32.4 Calculate the percentage increasThus, the correct answer is option A.